OpenStudy (anonymous):

Consider the 2-dimensional force field F = (4e^(-2x)+3y^3) i +9xy^2 j a) is F conservative? find a potential function f(x,y) whose gradient is F. b) find the work done by the force field F in moving an object from P(0,1) to Q(1,2) along the path y= sin( x pi/2) from x= 0 to x= 1.

OpenStudy (uri):

Do you know what it means for F to be conservative?

OpenStudy (anonymous):

ok I got part a) F is conservative because \[\frac{ dP }{ dy } = \frac{ dQ }{ dx } = 9y^2\]

OpenStudy (anonymous):

yeah, but I don't know what to do in part b)

OpenStudy (anonymous):

the potential function is \[3xy^3 -2e ^{-2x} \]

OpenStudy (anonymous):

right? @uri

OpenStudy (uri):

So the idea of F being conservative really means that the work done in a conservative field is the same regardless of the path taken.

OpenStudy (uri):

So, we can take a simpler path if we want to.

OpenStudy (uri):

Let me check

OpenStudy (anonymous):

so we don't have to use \[y = 1+\sin (\frac{ \pi x }{ 2 }) \]?

OpenStudy (mathmate):

Py=Qx => conservative is good. For part b, you need to find f such that \(\Delta f=F \) Can you do that?

OpenStudy (mathmate):

Sorry, it's not Delta f, it's grad. the triangle should be upside down. \(<f_x(x,y),f_y(x,y)> = F(x,y)\)

OpenStudy (anonymous):

I'm not sure @mathgirl2012 from part a we know that \[Δf=F\]

OpenStudy (uri):

You're right @Ldaniel

OpenStudy (anonymous):

so what's \[f\] the path given in part b? @mathmate

OpenStudy (uri):

Correct. We could use y = x+1 if we want to because it is simpler and the field is conservative

OpenStudy (mathmate):

Sorry, this is the second part of (a), not (b). This is what you need to find, from the given P and Q of F, and f has to satisfy the conditions I gave.

OpenStudy (uri):

The path given in part b is one possible path, but since F is conservative then the path taken is irrelevant. The Force is the same (thus the answer to the question) if you use the path y = 1 +sin(pix/2) or y = 1+x.

OpenStudy (mathmate):

If the field is conservative, work done is just f(b)-f(a).

OpenStudy (mathmate):

Like @uri said, independent of path.

OpenStudy (mathmate):

But you need to find f(x,y) first.

OpenStudy (anonymous):

ok so what f(x,y) are we talking about here?

OpenStudy (anonymous):

so Q(1,2) - P(0,1) using y = 1+x?

OpenStudy (mathmate):

Not really, first find f(x,y) from P, Q that satsfy \(<f_x(x,y),f_y(x,y)> = F(x,y)=<P(x,y),Q(x,y)>\)

OpenStudy (mathmate):

Then if we need work done from a to b, then work = f(b)-f(a).

OpenStudy (anonymous):

what surface do i use to find \[fx(x,y),fy(x,y)\]

OpenStudy (mathmate):

@Ldaniel Your proposed potential function is close. Can you check if it satisfies the above condition? If not, then you can make adjustments.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

ok so for \[f_x (x,y) = 3y^3 +4 e ^{-2x}\]

OpenStudy (anonymous):

and \[f_(x,y) = 9y^2x\]

OpenStudy (mathmate):

Can you check if \(F = (4e^{2x}+3y^3)i +9xy^2 j \) or \(F = (4e^{-2x}+3y^3)i +9xy^2j \)

OpenStudy (anonymous):

i'm getting the second one

OpenStudy (mathmate):

But is the one given in the question the correct one?

OpenStudy (anonymous):

I edit the question it forgot the negative sign sorry

OpenStudy (mathmate):

If the first one that you need, then you need to make a slight adjustment to f(x,y).

OpenStudy (mathmate):

Ok, so your potential function is correct, because if you derive wrt x, you get P, and wrt y, you get Q, right?

OpenStudy (anonymous):

yeah so not what do i do?

OpenStudy (mathmate):

Almost done!

OpenStudy (mathmate):

Work done from A(0,1) to B(1,2). I want to avoid P,Q which are confusing.

OpenStudy (mathmate):

Can you find that? It is independent of path, as @uri said.

OpenStudy (anonymous):

to what fuction do i plot those points, to the path?

OpenStudy (mathmate):

W=f(B)-f(A)

OpenStudy (anonymous):

I use the potential function?

OpenStudy (mathmate):

yes. \( \int gradf.dr = f(r(b)-f(r(a)) \) That is why is so much easier working with a conservative field.

OpenStudy (anonymous):

i'm getting 25.729... does that look right?

OpenStudy (mathmate):

I don't usually evaluate powers of e. Leave them in your answer, see what you get.

OpenStudy (anonymous):

(24-2e^(-2))- (-2)

OpenStudy (anonymous):

26 - 2e^(-2)

OpenStudy (mathmate):

Yes, that's what I got too! Good job!

OpenStudy (anonymous):

thank you so much for your help :)

OpenStudy (mathmate):

You're welcome! :) Just for curiosity, how did you get f(x,y) at the very beginning?

OpenStudy (anonymous):

integration

OpenStudy (mathmate):

Good, integrate P wrt x?

OpenStudy (anonymous):

\[\int\limits\limits4e ^{-2x}+ 3y^3 dx\]

OpenStudy (anonymous):

yeah, well I did both P and Q

OpenStudy (mathmate):

Did you differentiate with respect to y to make sure you get Q? ok that's good!

OpenStudy (anonymous):

yeah

OpenStudy (mathmate):

Very well! congrats!

OpenStudy (anonymous):

and i got 3xy^3

OpenStudy (mathmate):

Yes, that is Q, so you made sure the condition grad f = F is satisfied.

OpenStudy (anonymous):

so thats why I know that integrate P wrt x was the potential function

OpenStudy (mathmate):

Perfect, so the only (easy) part missing was f(B)-f(A).

OpenStudy (mathmate):

Which you just did.

OpenStudy (anonymous):

yeah I was confuse because of the given path

OpenStudy (anonymous):

I didn't know what to do with it, but now i know to do nothing haha

OpenStudy (mathmate):

The path was necessary IF the field was not potential. That could be confusing sometimes.

OpenStudy (anonymous):

yeah

OpenStudy (mathmate):

I feel a little guilty because I seem to have pushed @uri away. Sorry about that.

OpenStudy (mathmate):

Anyway, good luck with your other ventures!

OpenStudy (anonymous):

thanks :)

OpenStudy (uri):

lol,It's cool. :)

OpenStudy (mathmate):

You're welcome! :)

OpenStudy (mathmate):

@uri you're really nice!

OpenStudy (anonymous):

thank you @uri :)